OpenStudy (moazzam07):

Which fraction is equivalent to negative 2 over 3? 2 over negative 3 negative 3 over 2 negative 2 over negative 3 2 over 3

OpenStudy (anonymous):

consider that a fraction is simply a division problem. In this case -2 divided by 3.

OpenStudy (anonymous):

so looking at your answer choices, which fraction, or division problem, would give you the same answer?

OpenStudy (tkhunny):

It is important to note that a fraction has THREE signs. Change any TWO and you don't change the value. \(\dfrac{-2}{3} = -\dfrac{2}{3} = -\dfrac{-2}{-3} = \dfrac{2}{-3}\)

OpenStudy (moazzam07):

a

OpenStudy (anonymous):

uh positive 2 over 3

OpenStudy (moazzam07):

oh so its D?

OpenStudy (anonymous):

@moazzam07 , not d. Look at the examples @tkhunny gave you.

OpenStudy (tkhunny):

Don't forget to look at the sign out in front of the fraction. That one counts, too.

OpenStudy (moazzam07):

Um 2 over -3